OpenStudy (wild_child117):

Solve for c. a=11+4b−4c c=a+4b+11−4 c=a−b+11−4 c=a+b−11−4 c=a−4b−11−4

OpenStudy (wild_child117):

@quickstudent

OpenStudy (wild_child117):

@zepdrix

OpenStudy (quickstudent):

Hmm I haven't seen choices like those before.

alones (alones):

So solve for C?

OpenStudy (wild_child117):

yes i need to get a good grade on this test to get a d instead of a f and im not good at math

563blackghost (563blackghost):

You need to get all the variables to one side to get an equation for c.... \(\huge{a=11+4b-4c}\) First we would subtract by 11 to get it to the other side.. \(\Large{a-11=(11-11)+4b-4c \rightarrow a-11=4b-4c}\) What would we do next?

OpenStudy (1davey29):

Correct me if I'm wrong, but aren't none of these right?

OpenStudy (wild_child117):

idk

563blackghost (563blackghost):

oops didnt see that you said it was a test.....sorry but i cant help

alones (alones):

Yeah we're not suppose to help on exams, and test..use your knowledge.. We're here to help you study and assist you (: